# Large Language Model Strategic Reasoning Evaluation through Behavioral Game Theory

### v2

@@ Abstract
Strategic decision-making involves interactive reasoning where agents adapt their choices in response to others, yet existing evaluations of large language models (LLMs) often emphasize Nash Equilibrium (NE) approximation, overlooking the mechanisms driving their strategic choices. To bridge this gap, we introduce an evaluation framework grounded in behavioral game theory, disentangling reasoning capability from contextual effects. Testing 22 state-of-the-art LLMs, we find that GPT-o3-mini, GPT-o1, and DeepSeek-R1 dominate most games yet also demonstrate that the model scale alone does not determine performance. In terms of prompting enhancement, Chain-of-Thought (CoT) prompting is not universally effective, as it increases strategic reasoning only for models at certain levels while providing limited gains elsewhere. Additionally, we investigate the impact of encoded demographic features on the models, observing that certain assignments impact the decision-making pattern. For instance, GPT-4o shows stronger strategic reasoning with female traits than males, while Gemma assigns higher reasoning levels to heterosexual identities compared to other sexual orientations, indicating inherent biases. These findings underscore the need for ethical standards and contextual alignment to balance improved reasoning with fairness.
Keyword: Multi-agent, LLM, Reasoning, Fairness, Game Theory
@@ 1 Introduction
The rapid development of large language models (LLMs) and generative AI has significantly broadened their applications, transitioning from basic text generation [54] and completion tasks to serving as sophisticated agents [59, 76, 29, 25]. While benchmarks for language and math tasks assess specific abilities of LLMs [45, 31, 75], practical deployment alongside other agents or humans presents a broader and more complex challenge. This necessitates evaluating reasoning capabilities that integrate and extend these individual skills on language, vision, or math [39, 38, 42, 53, 66].
When AI agents are deployed in real-world applications, they often encounter decision-making scenarios requiring cooperation or competition with other entities—a process known as strategic reasoning [80, 43, 30]. One-shot strategic reasoning specifically refers to an agent’s ability to make a single, one-time action based on its reasoning in an interactive multi-agent environment. Consider an AI agent participating in a high-stakes negotiation, such as allocating resources in a disaster relief effort among multiple parties. The agent must evaluate the needs and strategies of other agents, analyze available resources, and understand the broader context to make a single, impactful decision. This one-shot reasoning is critical, as the decision cannot be altered once made, and it directly affects the outcomes for all participants. Such scenarios highlight the importance of robust frameworks for evaluating and improving one-shot strategic reasoning in AI systems [56, 20].
@@
Previous research on LLMs’ strategic reasoning has predominantly focused on their ability to achieve Nash Equilibrium (NE) in game-theoretic settings [80, 74, 30, 47]. While NE is a cornerstone concept in evaluating rational behavior, relying solely on whether LLMs can reach NE provides an incomplete assessment of their reasoning capabilities [48, 44]. The primary limitation of using NE as a measurement lies in its strict assumptions and focus on outcomes rather than the decision-making processes [3]. NE assumes perfect rationality, which fails to account for the variability and bounded reasoning capability inherent in real-world problems. For LLMs—probabilistic models trained on human-generated data—this assumption is particularly problematic [79, 55]. Their stochastic nature makes NE impractical as a comprehensive evaluation metric, and its assumption of perfect rationality falls into the circular reasoning fallacy when assessing LLMs’ rationality.
In this work, we move beyond traditional game theory and instead explore the factors shaping their strategic behavior. Drawing on the Truncated Quantal Response Equilibrium (TQRE) from behavioral game theory [67], we identify two key causes of deviations from optimal responses: (1) reasoning capability, reflecting an agent’s ability and depth to think rationally and adaptively, and (2) contextual structure, which influences reasoning through environmental complexity and acts as a confounder. Based on these factors, we develop a mechanism that controls for contextual effects and evaluates LLMs’ strategic reasoning across multi-level contexts and games. In summary, our key contributions are as follows:
@@
    We propose an evaluation framework to systematically assess reasoning capability while accounting for contextual complexity. Testing 22 state-of-the-art (SOTA) LLMs on 13 abstracted real-world games using one-shot prompting. We find that while massive models like GPT-o1, GPT-o3-mini, and DeepSeek-R1 lead in most tasks, smaller models can outperform or match their reasoning in others. This highlights the potential of task-specific usability over sheer model size and architecture.
@@
    We evaluate the impact of Chain-of-Thought (CoT) prompting on reasoning across all models and find that it is not a universal enhancer. CoT promotes reasoning when a model’s capability is limited. However, it can introduce distractions for LLMs with higher reasoning levels, leading to suboptimal choices. This finding aligns with recent studies showing that while CoT improves mathematical and sequential reasoning, it can also amplify distractions, varying its overall effectiveness.
@@
    Lastly, embedding demographic features before reasoning revealed biases even in advanced models like DeepSeek-R1 or GPT-4o when processing minority features. In addition, distilled knowledge could also cause bias injected into previous non-biased models such as DeepSeek-R1-Distilled-LLaMA. This suggests that superior reasoning ability does not necessarily yield desirable or ethical outcomes, highlighting the need for a balanced approach and calibration in future LLM development.
2 Background and Related Literature
2.1 LLM Behavior Study
Individual Decision-making Recent research [19, 80, 74] has increasingly focused on investigating the behavior of LLMs as both individual decision-makers and strategic decision-makers, exploring their alignment with human behavior [5, 15, 46, 24]. From an economic behavior perspective, LLMs demonstrate patterns resembling human behaviors, such as preferences under uncertainty and risk aversion [52]. From a social science perspective, LLMs demonstrate human-like behavior in moral judgment experiments, exhibiting reasoning patterns similar to those of humans [23, 51, 68].
In the context of individual decision-making, several studies [52, 60, 46] have revealed notable parallels between LLMs and humans, including differences in risk preferences, probability weighting, and loss aversion. Others [40, 52] highlight decision-making inconsistencies in LLMs when profile information with minority attributes is embedded, where LLMs may demonstrate contradictory interpretations of group-specific behaviors.
Strategic Decision-making Recent research also evaluates LLMs in strategic contexts, where reasoning involves dynamic, adaptive responses in cooperative or competitive environments. For instance, several studies explore LLMs’ abilities in achieving rationality within game-theoretic frameworks, such as mixed-strategy NE games [69] and classic strategic games like the prisoner’s dilemma [6, 37]. Some works investigate the adaptability of LLMs to various prompts that influence cooperative and competitive behavior [64], and others evaluate their capacity to replicate human-like reasoning patterns in multi-agent settings [1, 77]. While these studies provide an important starting point, many are limited to binary assessments of whether LLMs meet NE [47, 26] without exploring the underlying mechanisms driving their strategic reasoning capabilities. This creates a challenge in explaining deviations from optimal strategies and introduces circular reasoning in LLM evaluation, where the assumption of perfect rationality undermines the very purpose of evaluating reasoning processes. Notably, [27] and [81] help relieve these issues by exploring reasoning through stepwise evaluation and Cognitive Hierarchy (CH) models, respectively, but remain constrained by deterministic context assumptions.
2.2 Behavioral Game Theory
Behavioral game theory models strategic interactions by incorporating bounded rationality, decision noise, and social preferences, addressing deviations from Nash Equilibrium observed in real-world and experimental games [8, 11].
One key finding in behavioral game theory is the prevalence of stochastic decision-making, where players respond probabilistically to payoffs rather than always choosing optimally [41, 14]. For example, in private-value auctions, where bidders have independent valuations, overbidding is common due to bounded rationality and noise [33]. In public goods games, where individuals decide how much to contribute to a shared resource, contributions initially exceed Nash Equilibrium (NE) predictions due to fairness before declining with payoff adaptation [28]. These behaviors are effectively modeled by Quantal Response Equilibrium (QRE), which incorporates decision noise to explain such deviations [57].
Another key insight is the heterogeneity in reasoning levels, where players vary from relying on simple heuristics to engaging in iterative strategic thinking [16, 12, 4]. Cognitive hierarchy models capture this variation, showing that limited reasoning often dominates in complex environments [12, 70].
Structured environments allow limited reasoning to yield near-optimal payoffs, while complex settings amplify stochastic deviations [58]. For instance, in real-world experiments on bargaining games, where players take turns making and responding to offers, players deviate further from Nash Equilibrium (NE) as complexity increases [34]. Truncated Quantal Response Equilibrium (TQRE) [67] extends prior probabilistic models by integrating bounded rationality and reasoning depth, capturing the interplay between environmental structure, decision noise, and strategic behavior.
Behavioral game theory offers superior predictive power and richer insights into strategic interactions by incorporating bounded rationality, decision noise, and heterogeneous reasoning, addressing complexities that traditional game theory cannot fully capture [18, 12, 9, 10]. Building on TQRE, our work is the first to focus on eliciting and evaluating reasoning capabilities in LLMs while carefully accounting for contextual confounders in multi-agent environments.
3 Preliminary
Strategic decision-making is fundamentally grounded in game theory, which provides a framework for analyzing interactions between rational agents. It evolves into behavioral game theory by incorporating bounded rationality and cognitive limits. The TQRE model, which integrates stochastic decision-making and bounded reasoning, is particularly suited for evaluating LLMs as it aligns with their probabilistic nature and limited cognitive depth.
In a strategic-form game $\Gamma=(N,\{A_{i}\}_{i=1}^{N},\{u_{i}\}_{i=1}^{N})$, players choose strategies from their respective action sets to maximize their utility functions. A strategy profile $(\sigma_{1},\ldots,\sigma_{N})$ achieves a NE if no player can unilaterally improve their expected utility. Formally, for all players $i\in N$: $U_{i}(\sigma_{i}^{*},\sigma_{-i}^{*})\geq U_{i}(\sigma_{i},\sigma_{-i}^{*}),\quad\forall\sigma_{i}\in\Delta(A_{i})$, where $\sigma_{i}$ represents a mixed strategy over the action set $A_{i}$, and $\sigma_{-i}$ denotes the strategies of all other players.
The expected utility (payoff values) for player $i$ under $(\sigma_{i},\sigma_{-i})$ depends on both their own strategy and the strategies of others:
$U_{i}(\sigma_{i},\sigma_{-i})=\sum_{a\in A}\sigma_{i}(a_{i})\sigma_{-i}(a_{-i})u_{i}(a).$
Although NE provides a sound theoretical benchmark and is useful in computational tasks, its assumptions of infinite reasoning capacity and deterministic behavior often fail to align with empirical observations [10, 65, 62, 28, 17].
Quantal Response Equilibrium (QRE) introduces stochastic decision-making into the framework to address these limitations. In QRE, players choose actions with probabilities proportional to the expected utilities of the actions. The probability of player $i$ selecting action $a_{ij}$ is defined as:
$p_{ij}(\lambda,U)=\frac{\exp(\lambda U_{ij})}{\sum_{a\in A_{i}}\exp(\lambda U_{ia})},$
where $U_{ij}$ is the expected utility of action $a_{ij}$, and $\lambda>0$ controls decision precision. As $\lambda\to\infty$, QRE converges to NE, selecting the highest-utility action deterministically, while a lower $\lambda$ introduces more randomness.
TQRE extends QRE by introducing two key refinements: heterogeneity in reasoning levels and contextual scaling of payoff sensitivity. TQRE uses the Cognitive Hierarchy framework to model heterogeneity in reasoning, where players differ in strategic depth. Reasoning levels $k$ follow a truncated Poisson distribution: $f_{k}=\frac{\tau^{k}e^{-\tau}}{k!}$ where $\tau$ is the mean reasoning level. A level-$k$ player computes expected utility by aggregating strategies of lower-level players ($h<k$). For action $a_{ij}$, their expected utility is:
$U_{ij}^{(k)}=\sum_{a_{-i}\in A_{-i}}\prod_{h=0}^{k-1}\sigma_{-i}^{(h)}(a_{-i})u_{i}(a_{ij},a_{-i}),$
[…]
The probabilistic choice for a level-$k$ player is given by:
$p_{ij}^{(k)}=\frac{\exp(\lambda_{k}U_{ij}^{(k)})}{\sum_{a\in A_{i}}\exp(\lambda_{k}U_{ia}^{(k)})}.$
To compute the overall probability of choosing action $a_{ij}$ in the population, TQRE aggregates over all reasoning levels:
$p_{ij}=\sum_{k=0}^{\infty}f_{k}\cdot p_{ij}^{(k)},$
$f_{k}$ is the probability of a player with reasoning level $k$.
To summarize, TQRE framework introduces two key parameters: $\tau$, the reasoning level of the sample, and $\gamma$, the contextual scaling. These parameters jointly determine the rationality parameter $\lambda$ for the reasoning levels across the sample, balancing cognitive capability, represented by $\tau$, and game complexity, represented by $\gamma$. As $\tau\to\infty$ and $\gamma\to\infty$, the model converges to NE, where all players are perfectly rational and deterministic.
4 Framework and Design
Figure 1: Three main steps in the framework include (1) converting decision-making challenges to abstracted games, (2) prompting agents to make decisions, and (3) applying the TQRE model to estimate parameters of interest, assess demographic bias, and evaluate prompt effects.
The framework for evaluating LLMs’ strategic reasoning is outlined in Figure 1. The framework accounts for the end-to-end steps from game abstraction and designing for behavioral evaluation of the LLMs.
@@ 4.1 Game Library Design
The game library consists of diverse game types and payoff matrices. Each game incorporates different context levels by design to enable precise estimation of reasoning levels and dynamically assess the models’ reasoning boundaries. Drawing on established frameworks in behavioral game theory, we construct games under two primary categories: complete information games, where payoffs and strategies are common knowledge, and incomplete information games, where players form probabilistic beliefs about unknown elements of the game.
The complete information games include four distinct subtypes designed to capture a range of interactions:
@@
    Competitive (zero-sum) games: Players’ interests are strictly opposed, with one’s gain equaling the other’s loss, as the total payoff remains constant. These games require adversarial reasoning to determine optimal strategy.
@@
    Cooperation games (e.g., Stag Hunt): These games emphasize the trade-off between mutual coordination and individual incentives. Players achieve the highest payoffs when they cooperate, but the risk of unilateral defection can lead to suboptimal outcomes. Success depends on trust and assurance in the other player’s actions.
@@
    Mixed-motive games (e.g., Prisoner’s Dilemma): Characterized by a tension between one’s incentives to defect and potential gains from mutual cooperation. These games are essential for studying trust, reciprocity, and strategic foresight.
@@
    Sequential games: One player (leader) moves first, while the other (follower) observes and responds optimally. These games involve backward induction, where the follower’s optimal strategy influences the leader’s initial decision.
The incomplete information games simulate environments in which players must infer unknown elements, such as their opponent’s type or payoff structure. These games incorporate probabilistic reasoning, with opponent types $t$ distributed according to $\mathbf{P}(t)$. Two key games are:
@@
    Bayesian coordination games: Players make decisions based on posterior beliefs about their opponent’s type, adjusting their strategies accordingly. The uncertainty about the opponent’s characteristics influences equilibrium selection and coordination success.

### v3

# LLM Strategic Reasoning: Agentic Study through Behavioral Game Theory

@@ Abstract
What does it truly mean for a language model to “reason” strategically, and can scaling up alone guarantee intelligent, context-aware decisions? Strategic decision-making requires adaptive reasoning, where agents anticipate and respond to others’ actions under uncertainty. Yet, most evaluations of large language models (LLMs) for strategic decision-making often rely heavily on Nash Equilibrium (NE) benchmarks, overlook reasoning depth, and fail to reveal the mechanisms behind model behavior. To address this gap, we introduce a behavioral game-theoretic evaluation framework that disentangles intrinsic reasoning from contextual influence. Using this framework, we evaluate 22 state-of-the-art LLMs across diverse strategic scenarios. We find models like GPT-o3-mini, GPT-o1, and DeepSeek-R1 lead in reasoning depth. Through thinking chain analysis, we identify distinct reasoning styles—such as maximin or belief-based strategies—and show that longer reasoning chains do not consistently yield better decisions. Furthermore, embedding demographic personas reveals context-sensitive shifts: some models (e.g., GPT-4o, Claude-3-Opus) improve when assigned female identities, while others (e.g., Gemini 2.0) show diminished reasoning under minority sexuality personas. These findings underscore that technical sophistication alone is insufficient; alignment with ethical standards, human expectations, and situational nuance is essential for the responsible deployment of LLMs in interactive settings.
@@ 1 Introduction
The rapid development of large language models (LLMs) and generative AI has significantly broadened their applications, transitioning from basic text generation [41] and completion tasks to serving as sophisticated agents [45, 60, 21, 17]. While existing benchmarks primarily assess LLMs on isolated language or math tasks [33, 23, 58], real-world deployment often demands more complex forms of decision-making, particularly strategic reasoning, where agents must interact with other entities whose actions directly influence the outcomes [64, 31, 22]. Specifically, one-shot strategic reasoning refers to an agent’s ability to select a single, irreversible action where the outcome depends on both its own choice and that of other agents in this interaction. Consider an AI agent operating in a business environment where strategic interaction is critical. In procurement, for instance, it must anticipate supplier counteroffers when negotiating orders; in advertisement bidding, it predicts competitors’ bids to set optimal prices. In each case, the agent faces a one-shot decision that directly impacts its payoff and cannot be revised post-execution [43, 12].
Related Literature and Research Gap. Research on LLM decision-making typically begins by examining their behavior as independent decision-makers [4, 11, 34, 16]. From an economic perspective, LLMs exhibit human-like patterns in preferences under uncertainty and risk [40]. In social science, they have demonstrated alignment with human responses in moral judgment and fairness-related tasks [15, 39, 52]. Other studies further reveal alignments between LLMs and human judgments across diverse decision contexts [34, 46, 40]. However, when demographic profiles are embedded into prompts, systematic inconsistencies and fairness concerns emerge [29, 40].
@@
Strategic decision-making extends individual reasoning into interactive contexts, where outcomes depend not only on one’s own choices but also on the anticipated actions of others. Existing evaluations of LLM strategic reasoning are predominantly grounded in game theory. Several studies explore LLMs’ abilities in achieving rationality within game-theoretic frameworks, such as mixed-strategy Nash Equilibrium (NE) games [53] and classic strategic games like the prisoner’s dilemma [5, 28]. Some works investigate the adaptability of LLMs to various prompts that influence cooperative and competitive behavior [50], and others evaluate their capacity to replicate human-like reasoning patterns in multi-agent settings [1, 61]. While these studies provide a good starting point, many are limited to binary assessments of whether or not LLMs meet NE [35, 18] without quantifying their strategic reasoning capabilities or exploring the underlying mechanisms. While NE is a cornerstone concept in evaluating rational behavior, evaluating solely whether LLMs can achieve NE is an incomplete assessment of their reasoning capabilities [36, 32]. The primary limitation of using NE as a measurement lies in its strict assumptions and focus on outcomes rather than the decision-making processes [3]. It makes it difficult to interpret why LLMs deviate from optimal strategies. Also, NE assumes perfect rationality, which fails to account for the variability and bounded reasoning capability inherent in real-world problems. For LLMs—probabilistic models trained on human-generated data—this assumption is particularly problematic [63, 42]. Their stochastic nature makes NE impractical as a comprehensive evaluation metric, and its assumption of perfect rationality falls into the circular reasoning fallacy when assessing LLMs’ rationality. Recent work begins to address these issues by incorporating stepwise evaluation [19] or applying models like Cognitive Hierarchy (CH) [65], but these approaches still rely on deterministic assumptions about context and agent type.
Figure 1: Framework overview
Key Contributions. To bridge this gap, we propose a framework, as shown in Figure 1, that moves beyond assessing LLMs with traditional game theory. We develop a method to quantify and characterize LLMs’ strategic decision-making, capturing bounded rationality, response to incentives, and belief formation. Our approach combines a diverse library of matrix games with a structured modeling framework based on Truncated Quantal Response Equilibrium (TQRE) stemming from behavioral game theory [51, 59]. The key contributions are as follows:
@@
    We introduce an evaluation framework for assessing strategic reasoning in LLMs under various conditions. Applying to 22 state-of-the-art (SOTA) models on 13 abstracted real-world games, we find that massive models such as GPT-o3-mini and DeepSeek-R1 achieve top reasoning depth across most tasks, while smaller models occasionally match or even outperform them in specific game types.
@@
    Based on our evaluation results, we investigate why different models exhibit varying reasoning depth across games by analyzing the reasoning chains of three top-performing models in baseline competitive, cooperative, and mixed-motive games. Analysis reveals that each model’s performance is closely tied to its dominant reasoning style, but does not benefit well from long reasoning chains.
@@
    Finally, we examine the social effects and alignment of LLMs relative to human behavior when demographic features are assigned. We find that embedding demographic attributes prior to reasoning can reveal biases, even in advanced models such as DeepSeek-R1 and GPT-4o, particularly when processing minority-related features. This indicates that superior reasoning capabilities do not inherently lead to desirable or ethical outcomes, underscoring the need for careful calibration and a balanced approach in future LLM development.
2 Theoretical Foundations
2.1 Literature on Behavioral Game Theory
Classical game theory begins with the formalism of NE, in which subjects are assumed to be fully rational: they possess unlimited reasoning capacity, have perfect knowledge of the game, and make deterministic best responses [48]. While mathematically sound, empirical evidence demonstrates frequent deviations from NE. For example, subjects often exhibit bounded rationality and stochastic behavior, such as overbidding in auctions or early cooperation in public goods games [25, 20]. This gap between theory and observed behavior prompted the emergence of behavioral game theory, which progressively relaxes NE’s strong assumptions [7, 9, 8, 54].
The first core insight is that subjects do not always select the utility-maximizing option deterministically. Instead, they respond to incentives in a probabilistic pattern, leading to stochastic choice models, as proposed in the Quantal Response Equilibrium (QRE) model [44, 30]. Another observation is that subjects differ in their depth of strategic reasoning levels. Rather than assuming that all players engage in unbounded thinking, models such as CH posit that individuals operate at varying levels of sophistication [54, 9]. To capture both bounded reasoning and probabilistic behavior, we adopt the TQRE model, which combines the key behavioral refinements of QRE and CH. Instead of assuming convergence to a Nash fixed point, TQRE models subjects as reasoning at limited depths and responding stochastically, thus capturing more realistic strategic behavior, structured as below.
2.2 Modeling Preliminaries
Bounded‐rational Belief Formation.
Each subject draws a reasoning level $k\sim\mathrm{Poisson}(\tau),\quad\tau>0,$ where $\tau$ is the average depth of strategic thought. A level-$k$ subject does not know each opponent’s exact level $h<k$, so it mixes their level-$h$ strategy distributions $\sigma_{j}^{(h)}:A_{j}\to[0,1],\;\sum_{a_{j}}\sigma_{j}^{(h)}(a_{j})=1,$ using weights $W_{h}^{(k)}$, to form the weighted average of lower-level strategies. The marginal belief is
$\bar{\sigma}_{j}^{(k)}(a_{j})=\sum_{h=0}^{k-1}W_{h}^{(k)}\,\sigma_{j}^{(h)}(a_{j}).$
Given that opponents act independently, the joint belief over all others’ actions $a_{-i}\in A_{-i}$ is
$\Pr(a_{-i})=\prod_{j\neq i}\bar{\sigma}_{j}^{(k)}(a_{j})$
, where the notation $a_{-i}\$stands for the actions of all players except $i$.
Expected Utility (EU) and Stochastic Choice.
Given $\Pr(a_{-i})$, a level-$k$ subject’s expected payoff for each action $a_{ij}\in A_{i}$ is
$U_{ij}^{(k)}=\sum_{a_{-i}\in A_{-i}}\Pr(a_{-i})\,u_{i}(a_{ij},a_{-i})$
, where $u_{i}$ is the payoff function. To allow for decision noise, set $\lambda_{k}=\gamma\,k$ with $\gamma>0$, and convert utilities into choice probabilities:
$p_{ij}^{(k)}=\frac{\exp\!\bigl(\lambda_{k}\,U_{ij}^{(k)}\bigr)}{\sum_{a\in A_{i}}\exp\!\bigl(\lambda_{k}\,U_{ia}^{(k)}\bigr)}.$
Population‐Level Aggregation.
Finally, the overall probability that a randomly selected subject plays $a_{ij}$ is the mixture over reasoning levels:
$p_{ij}=\sum_{k=0}f_{k}\;p_{ij}^{(k)},\quad f_{k}=\frac{\tau^{k}\,e^{-\tau}}{k!},$
By estimating $\tau$, we measure the reasoning depth LLMs entertain when making an interactive decision. We thus recover a cognitively grounded metric of bounded strategic reasoning that moves beyond NE-based assumptions to characterize how LLMs reason in multi-agent environments. For more details of the model specification, see the Appendix A.
Our modeling approach offers two key advantages over traditional NE-based evaluations. First, by incorporating bounded rationality and heterogeneous reasoning levels, it avoids the circularity of assuming optimal behavior to test for rationality. Instead, it enables empirical estimation of strategic depth without presupposing NE play. Second, it captures the anticipatory nature of strategic reasoning: each agent’s EU is grounded in beliefs about opponents’ actual behavior, where $\tau$ quantifies the model’s depth of recursive belief modeling, i.e., layers of back-and-forth reasoning about others’ reasoning it engages in. This provides a cognitively meaningful measure of interactive sophistication.
Interpreting $\gamma$ (decision precision).
In the logit choice rule, $\gamma$ scales how deterministically utilities are translated into actions conditional on beliefs. Larger $\gamma$ yields more deterministic (near–best-response) behavior, while lower $\gamma$ reflects noisier or more exploratory choice behavior. Importantly, $\gamma$ does not encode preferences over outcomes—it captures the consistency of choice given utilities and beliefs. Because $\gamma$ can be context- and scale-dependent, we report it primarily to avoid conflating shallow reasoning with stochastic execution, while using $\tau$ as our core measure of strategic sophistication. Full $\gamma$ estimates are reported in Appendix B.
3 Experiment Setup and Estimation Framework
@@ 3.1 Game Library Design
We develop a library of 13 matrix games spanning 7 core types from behavioral game theory, grouped into complete-information (fully known payoffs) and incomplete-information (need to reason with uncertainty) settings. Each game varies in stakes and strategic structure.
Complete Information Games include four classic structures as the examples in Table 1(d). Each cell displays the payoffs for Player 1 (row chooser) and Player 2 (column chooser), respectively. Players choose their strategies, and the payoffs reflect the resulting outcome for each pair of chosen actions.
@@
    Competitive: Each player’s gain is the other’s loss, which tests adversarial reasoning.
@@
    Cooperation: Risky coordination, where mutual cooperation yields the highest joint payoff.
@@
    Mixed-Motive: Social dilemmas balancing private vs. joint outcomes (e.g., Prisoner’s Dilemma).
@@
    Sequential: temporal structure, with Player 1 moving first and Player 2 responding; we elicit decisions from the first mover.
Table 1: Representative matrices for complete information games
10,-10 0,5 -5,8
-10,10 5,0 8,-5
0,0 5,-5 -5,5
(a) Competitive
8,8 0,7
7,0 5,5
(b) Cooperation
3,3 0,5
5,0 1,1
(c) Mixed-Motive
0,5 0,3 0,0
5,2 3,3 -1,-1
2,4 4,3 0,-2
(d) Sequential
Incomplete Information Games incorporate belief-based reasoning about unknown payoff structures or player type. Examples of payoff matrices are shown in Table 2(d).
@@
    Bayesian Coordination: Players choose actions under uncertainty over the governing payoff matrix. Agents are given priors (e.g., 30% vs. 70%) over two types.